My question is about ratio.
The ratio of two numbers is 3:5. If 4 are both added to both numbers the resulting ratio will be 7:11, find the smaller number.
I hope you can help me

Answer by rothauserc(4718)

x/y = 3/5
5x = 3y
y = 5x/3
(x+4) / (y+4) = 7/11
11*(x+4) = 7 * (y+4)
11x +44 = 7y +28
11x = 7y -16
substitute for y in second ratio with value for y in first ratio
11x = 7*(5x/3) -16
11x = 35x/3 -16
multiply both sides of = by 3
33x = 35x -48
-2x = -48
x = 24
use first ratio to solve for y
y = (5*24) / 3 = 40
note that 24/40 = 3/5 and
28/44 = 7/11
the smallest number is x = 24